Calculate the energy associated with the magnetic field of a 213-turn solenoid in which a current of 1.77 A produces a magnetic flux of 3.83 time sign 10^-4 T.m^2 in each turn.  
Answer #1

here,


magnetic fluc , phi = 3.83 * 10^-4 T.m^2


Energy stored in inductor E = 0.5 * L * i^2 .. (i=current)

For long coil L = u0 * A*N^2/l ..

(l=length, A=Area, N=turns)

combining ..

E = 0.5 * (u0 * AN^2/l)i^2 = 0.5 * (u0 * N^2i^2)(A/l) ◄(Eq1)

getting (A/l) ..

B through coil given by B = u0 * Ni/l

Flux through coil B*A = A(u0 * Ni/l) = (A/l)u0 * Ni

From data, flux = {3.83 * 10^-4 *N}

combining ..

(A/l) = {3.83 * 10^-4 }/u0 * Ni = (3.83 * 10^-4) /(u0 N* i)◄(Eq2)

Putting it all together ... comb.Eq1&2

E = 0.5 * (u0 * N * i^2)(3.83 * 10^-4 /u0 * i*N)

E = 0.5 * (N * i)(3.83 * 10^-4)


number of turns , N = 213

current , I = 1.77 A

E = 0.5 * ( 213 * 1.77) * ( 3.83 * 10^-4)

E = 0.0722 J

the energy stored is 72.2 mJ